OpenStudy (anonymous):

simplify; |3n/n-3..-3|

OpenStudy (unklerhaukus):

\[\left|\frac{3n}{n-3}-3\right|\]?

OpenStudy (anonymous):

thats it

OpenStudy (unklerhaukus):

\[\left|\frac{3n}{n-3}-3\right|\]\[=\left|\frac{3n}{n-3}-3\times1\right|\]\[=\left|\frac{3n}{n-3}-3\times\frac{n-3}{n-3}\right|\]\[=\left|\frac{3n-3(n-3)}{n-3}\right|\]\[=\quad\dots\]

OpenStudy (anonymous):

please show how i cancel the top to obtain 9

OpenStudy (unklerhaukus):

\[=\left|\frac{3}{n-3}\right|\]\[=\left|\frac{3}{n-3}\times1\right|\]\[=\left|\frac{3}{n-3}\times\frac{n+3}{n+3}\right|\]\[=\left|\frac{3(n+3)}{n^2-9}\right|\]

OpenStudy (anonymous):

the correct answer is |9/n-3| in the book

OpenStudy (unklerhaukus):

ops -3 times -3 is 9, my last post is all wrong

OpenStudy (anonymous):

@UnkleRhaukus check your steps . \[\left| \frac{3n-3(n-3)}{n-3} \right|=\left| \frac{3n-3n+9)}{n-3} \right| =\left| \frac{9}{n-3} \right|\]

OpenStudy (unklerhaukus):

yeah i forgot my times tables for a moment there

OpenStudy (anonymous):

thank you very much shubham

OpenStudy (anonymous):

:)